The chickens at Colonel Thompson's Ranch have a mean weight of 1850 g, with a standard deviation of 160g. The weights of the chickens are closely approximated by a normal curve. Find the percent of all chickens having weights between 1418 g and 2156 g. Use the normal curve table

The percent of chickens with weights between 1418g and 2156 g is ____ %. ​(Type an integer or decimal rounded to two decimal places as​ needed.)
Answer #1

Answer:

Given,

Mean = 1850

Standard deviation = 160

P(1418 < X < 2156) = P((1418 - 1850)/160 < (x - mu)/s < (2156 - 1850)/160)

= P(-2.7 < z < 1.91)

= P(z < 1.91) - P(z < -2.7)

= 0.9719334 - 0.003467 [since from z table]

= 0.9685

P(1418 < X < 2156) = 96.85%